Darivation
V=√u^(2)+g^(2)t^(2)-2ugt sin(θ)
1
Expert's answer
2019-12-10T09:56:49-0500
"v=\sqrt{v_x^2+v_y^2}"

For projectile:

"v=\sqrt{(u\cos{\theta})^2+(u\sin{\theta}-gt)^2}"

"v=\sqrt{u^2((\cos{\theta})^2+(\sin{\theta})^2)+(gt)^2-2ugt \sin(θ)}"

"v=\sqrt{u^2+g^2t^2-2ugt \sin(θ)}"